A triangle has a side of length x. The second side of the triangle is 2 inches longer than the first side. The third side is 4 inches longer than the first side. The square of the first side plus the square of the second side is equal to the square of the third side. What is the length of each side of the triangle? Use the 5-step method.
The 5-step method is:
1. Select a variable to represent the basic unknown number.
2. Represent all the other unknown numbers in terms of that variable.
3. Write an equation that represents the relationships stated in the problem.
4. Solve the equation.
5. Check.

Answer by jorel1380(3719)

x2+(x+2)2=(x+4)2
x2+x2+4x+4=x2+8x+16
2x2+4x+4=x2+8x+16
x2-4x-12=0
(x-6)(x+2)=0
x=6 or -2
Throwing out the negative answer, we get the first side to be 6in., the second side to be 8in., and the third side to 10in..